I got this question in my homework and i didnt know how to answer it. Any help would be deeply appreciated.

*An intravenous solution contains 0.154 mol / L of sodium chloride. Calculate the mass of sodium in a 500 mL bag of this solution.
1
Expert's answer
2020-01-03T06:27:24-0500

Molarity of the given solution is given 0.154M

number of moles in 500 ml of the solution= M*V = 0.154*0.5 = 0.077 moles

number of moles of sodium will be same as moles of NaCl

hence moles of sodium= 0.77

molar mass of sodium =23

hence mass of sodium in the solution= molar mass *number of moles

= 23*0.077

=1.771grams